OpenStudy (anonymous):

Write each fraction in lowest terms: 4d/2d^6

OpenStudy (anonymous):

\[\frac{4d}{2d^6} = 2d^{1-6} = 2d^{-5} = \frac{1}{2d^5}\]

OpenStudy (jack1):

... 2 / 2d^5 ... 4d = 2 x 2d

OpenStudy (jack1):

@bloopman and @Zoobidoo